OpenStudy (anonymous):

please help!!!!! Brian has a 25-year adjustable rate mortgage that has a rate of 3.2% for the first 4 years. The amount of the mortgage is $179,000. What is the monthly payment during this initial period? $867.58 $596.67 $1,017.89 $615.76

OpenStudy (anonymous):

the intitial rate would be $615.76 per month

OpenStudy (amistre64):

P = Bk^(25*12) (1-k)/(1-k^(25*12)), k=1+.032/12

OpenStudy (anonymous):

my mind has just been blown i did it wrong, and i know why thank you @amistre64

OpenStudy (amistre64):

yep ....

OpenStudy (anonymous):

Wait so its not D ?

OpenStudy (amistre64):

why would it be D?

OpenStudy (anonymous):

take my medal away and give it to @amistre64

OpenStudy (anonymous):

I see it now, thank you so much!

OpenStudy (anonymous):

how did you know that formula @amistre64 ಠ_ಠ

OpenStudy (amistre64):

lol; i created it

OpenStudy (anonymous):

I posted another, can you guys help?

OpenStudy (amistre64):

i wrote out the recurrsion and simplified it; if i simplify it any more i get the textbooks formula ... but i find the textbook less intuitive

OpenStudy (anonymous):

Amistre64 should im still trying to figure out the formula

OpenStudy (anonymous):

ive never met anyone like you before, you are genuis

OpenStudy (amistre64):

let k be that compounding mess to clean this up; B is balance and P is payment \[B_o=B_0\] \[B_1=B_0k - P\] \[B_2=B_0k^2 - Pk -P\] \[B_3=B_0k^3 - Pk^2 -Pk -P\] this is a geometric sum of k on P \[B_n=B_ok^n-P\frac{1-k^n}{1-k}\] the nth Balance needs to be zero for the loan to be paid in full \[0=B_ok^n-P\frac{1-k^n}{1-k}\] solve for P \[B_ok^n=P\frac{1-k^n}{1-k}\] \[B_ok^n\frac{1-k}{1-k^n}=P\]

OpenStudy (anonymous):

@amistre64 is very smart and good at these problems!!

OpenStudy (anonymous):

I posted another if you can check it out

OpenStudy (amistre64):

i have my moments :) but im outta time for now, ill give a quick look at it tho

OpenStudy (anonymous):

that formula actually makes sense...mother of godಠ_ಠ

OpenStudy (amistre64):

lol :)

OpenStudy (amistre64):

yall have fun ... i might see wolverine this weekend